A student used the following data to calculate the mass % of total CO2 present in the soda. Be sure to use the current number of significant figures in your answer. Mass of soda bottle before opening = 533.4g. Mass of soda bottle after all CO2 was removed from the soda = 530.6 g.

a. 99.5% b. 0.52% c. 53% d. 99.48% e. 0.56%
Answer #1

Mass of soda bottle before opening = 533.4g

Mass of soda bottle after all CO2 was removed from the soda = 530.6 g

Mass of CO2 in bottle = Mass of soda bottle before opening - Mass of soda bottle after all CO2 was removed from the soda = 533.4g - 530.6 g = 2.8 g

Mass % of CO2 =(Mass of CO2 in bottle / Mass of soda bottle before opening)*100% = (2.8g/533.4g)*100% = 0.52%

option b